# Mpemba effect in a two-dimensional bistable potential

### 25 Mar 2026 v1

@@ Abstract
We present an exactly solvable model of the Mpemba effect in an overdamped Langevin system confined in a two-dimensional radially symmetric bistable potential. The potential is constructed as a piecewise quadratic-logarithmic function that is continuous and differentiable at the matching radii, enabling an exact mapping of the corresponding Fokker-Planck operator to a Schrödinger-type eigenvalue problem. The relaxation spectrum and eigenmodes are obtained analytically in each region in terms of confluent hypergeometric functions, with eigenvalues determined from matching conditions. Focusing on isotropic equilibrium initial states at inverse temperature $\beta_{\rm ini}$ quenched to a bath at inverse temperature $\beta$, we derive explicit expressions for the mode amplitudes governing long-time relaxation. We demonstrate that the coefficient of the slowest mode exhibits non-monotonic dependence on $\beta_{\rm ini}$ and identify a sufficient crossing condition for the Kullback-Leibler divergence in terms of the two slowest modes, if the global minimum of the potential is located far away from the origin and the second minimum exists near the origin. For corresponding parameters, we demonstrate that the Mpemba effect can be realized. Our results provide a rare example of an analytically tractable two-dimensional model exhibiting anomalous relaxation without any confining walls, extending previous one-dimensional constructions with a hard wall and clarifying the role of radial geometry in nonequilibrium relaxation phenomena.
@@ I Introduction
The Mpemba effect has been reported in a broad spectrum of classical systems, including colloidal particles in optical traps [6, 7], thermostated granular fluids [5, 8, 9, 10, 11, 12, 13], optical resonators [14, 15, 16], inertial suspensions [17, 18], spin glasses [19], molecular binary mixtures [20], phase transitions [21, 22], and a wide variety of Markovian model [4, 23, 24, 25, 26, 27, 28], non-Markovian models [29, 30] and active matters [31]. In addition, quantum analogues of the effect have been demonstrated in open quantum systems [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], establishing the Mpemba effect as a general nonequilibrium relaxation phenomenon.
@@
So far, we have only results for the motion of a particle in a discrete-level model, piece-wise linear potentials, and square box potentials [4, 5, 6, 57, 27]. There are no analytical studies of the motion of a particle confined to a smooth bistable potential, even in one-dimensional (1D) settings. Experimentally and theoretically, one-dimensional systems differ fundamentally from higher-dimensional systems, such as two-dimensional (2D) systems. Quite recently, Liu et al. [53, 54] have analyzed the motion in 1D asymmetric bistable potentials and found that the metastable picture is not correct for the realization of the Mpemba effect, but a confining wall is necessary to observe the Mpemba effect. To understand the robustness of the Mpemba effect, it is natural to analyze a 2D system to clarify the differences and common features in 1D and 2D systems, including whether confining walls are necessary to observe the Mpemba effect even in 2D.
@@
The contents of this paper are as follows. In the next section, we explain the model and setup, including the introduction of a solvable potential, and the mapping onto the Schrödinger equation. In Sec. III shows the mode-analysis of the mapped Schrödinger equation to obtain the solution of Schrödinger equation. In Sec. IV, we demonstrate the occurrence of the Mpemba effect by using the two-mode approximation. In Sec. V, we discuss our results. In Sec. VI, we present the summary of our results. In Appendix A, we briefly summarize the properties of hypergeometric functions we use. In Appendix B, we present the condition of the crossing of the KL divergence. In Appendix C, we discuss the contribution of higher-order terms of mode analysis. In Appendix D, we present the detailed conditions to have a peak of the slowest eigenmode against the initial temperature, mathematically.
@@ II.1 Formulation
Figure 1: A schematic of a bistable circular symmetric potential.
Let us consider the motion of a particle at a position $\bm{r}$ in a two-dimensional system. Here, the particle moves under the influence of a bistable potential $V(r)$ with $r:=|\bm{r}|$ (see Fig. 1). We assume that the motion of a particle obeys the dimensionless Langevin equation as
@@
with the bath temperature $T$.
@@
Note that $P(\bm{r},t)$ should be isotropic if we begin with an isotropic initial condition, i.e. $P(\bm{r},t)=P(r,t)$ because the dynamics is governed by a circular symmetric potential $V(r)$.
@@ II.2 A choice of solvable potential
In general, it is hard to obtain the exact solution of the Fokker-Planck equation with a double-well potential. However, if we choose a specific form of the potential $V(r)$, we can solve the Fokker-Planck equation. Thus, we aim to construct an exactly solvable radially symmetric double-well potential $V(r)$ in two dimensions. The requirements are: 1) $V(r)$ is continuous and differentiable across its domains, 2) $V(r)$ has a minimum at $r=0$, a maximum at $r=\xi$, and a second minimum at $r=\alpha>\xi$.
@@
where $r_{\pm}$ denote matching points, and $C_{\rm in},C_{\rm mid}$, and $C_{\rm out}$ are, respectively, constants to enforce continuity. As will be shown, the conditions $b_{\mathrm{mid}}\propto\xi^{2}$ and $b_{\mathrm{out}}\propto\alpha^{2}$ should be satisfied. Interestingly, this specific form leads to an exactly solvable Fokker-Planck equation.
The derivatives in each region lead to
@@
The second derivative of $V(r)$ leads to
@@
The potential $V(r)$ in Eq. (10) has two minima at $r=0$ and $r=\alpha$. Then, the values of $V(r)$ at the these minima are, respectively, given by $V(0)=0$, and
@@
where we have used Eqs. (9), (13), (14), and (15). We can consider the cases for (i) $V(\alpha)<0$, (ii) $V(\alpha)=0$, and (iii) $V(\alpha)>0$. If we choose $k_{\mathrm{in}}=k_{\mathrm{out}}=\xi=1$, the marginal line for $V(\alpha)=0$ is plotted as shown in Fig. 2.
@@ II.3.1 Framework
Demanding extrema for $V_{S}(r)$ at $r=0,\xi$, and $\alpha$ fixes
@@
We should note the following: i) Each segment of $V_{S}(r)$ is solvable quadratic plus inverse-square potential. ii) The solutions of the corresponding radial Schrödinger equation are expressible in terms of confluent hypergeometric functions. iii) The matching conditions at $r=r_{-}$ and $r=r_{+}$ enforce continuity of both $V(r)$ and $V^{\prime}(r)$, which automatically ensure the continuity of $V_{S}(r)$ and the physical wavefunctions. This construction provides a fully solvable model for a two-dimensional bistable potential with spherical symmetry.
@@
where we have used the relations of $\varphi_{m,n}(r)$ satisfying
@@
When we begin with an isotropic (equilibrium) condition Eq. (6), the solution of Schrödinger equation is independent of the second index $n$. Thus, the orthonormal condition the orthonormal relation
@@
$\langle\hat{A}\rangle(t)\approx 2\pi\int_{0}^{\infty}drrP_{\mathrm{eq}}(r,\beta)\hat{A}+2\pi\int_{0}^{\infty}drr\hat{A}\left[e^{-\beta V(r)/2}\{a_{2}\phi_{2}(r)e^{-\tilde{\lambda}_{2}t}+a_{3}\phi_{3}(r)e^{-\tilde{\lambda}_{3}t}\right],$ (34)
@@
Note that 1D model can be solved using a set of harmonic and anti-harmonic potentials [54]. The idea of using harmonic and anti-harmonic potentials to describe tunneling in an asymmetric double-well potential in quantum mechanics was introduced by Dekker [60] and later developed by Song [61, 62]. Our analysis shares similarities with previous papers [60, 61, 62], the target is completely different from them.
@@ II.3.2 Condition to observe the Mpemba effect
Let us consider Eqs. (II.3.1) and (34) in the large $t$ regime, where the relaxation term proportional to $a_{2}$ is only relevant. Then, the previous papers [4, 51, 53, 54] indicated that the condition to observe the Mpemba effect is that $a_{2}$ has a maximum or a minimum against $\beta_{\mathrm{ini}}$. Indeed, this condition is necessary because the peak of $a_{2}$ at $\beta_{\mathrm{ini}}^{*}$ corresponds to the slowest initial temperature to approach the final equilibrium state after the quench. This suggests that two relaxation processes, one starting from $\beta_{\mathrm{ini}}^{*}$ and the other from $\beta_{\mathrm{ini}}<\beta_{\mathrm{ini}}^{*}$, may cause a crossing of a thermodynamic observable. However, this condition may be insufficient, as it does not guarantee that the slowest eigenmode crosses another relaxation mode proportional to $a_{3}$. To improve the defect of the previous condition, only $a_{2}$, we will propose a more direct crossing condition for a monotone measure, such as the KL divergence, to observe the Mpemba effect later.
@@ II.3.3 Matching conditions
Since we adopt the connected potential Eq. (9), which can be converted into the effective potential Eq. (21), we should impose the connection conditions at $r=r_{-}$ and $r=r_{+}$. The connection conditions should read
@@
where $P_{\mathrm{in}}(r;t)$, $P_{\mathrm{mid}}(r;t)$, and $P_{\mathrm{out}}(r;t)$ are $P(r,t)$ for $0\leq r<r_{-}$, $r_{-}\leq r<r_{+}$, and $r\geq r_{+}$, respectively. Similarly, $J_{r}(r,\theta;t):=\bm{e}_{r}\cdot\bm{J}(r,\theta;t)$ is the radial component of the current $\bm{J}$ introduced in Eq. (5), and $J_{r,\mathrm{in}}$, $J_{r,\mathrm{mid}}$, and $J_{r,\mathrm{out}}$ are $J_{r}$ for $0\leq r<r_{-}$, $r_{-}\leq r<r_{+}$, and $r\geq r_{+}$, respectively. Since $U^{\prime}(r)$ is continuous at the connection points, the continuity condition of the current can read the continuity condition of $\partial P(r,t)/\partial r$ if the radial current $J_{r}$ is regular at the connection points. Using Eqs. (5) and (II.3.1), the continuity conditions can read
@@ III.1 Eigenvalue equation
Let us rewrite $\mathbb{H}$ introduced in Eq. (19) using a polar coordinate $r:=\sqrt{x^{2}+y^{2}}$ under circular symmetric cases as
@@
$\mu_{\mathrm{I},n}:=\frac{1+\nu_{\mathrm{I}}}{2}+\frac{\beta\left(\lambda_{n}-k_{\mathrm{I}}-\tfrac{k_{\mathrm{I}}|b_{\mathrm{I}}|}{2T}\right)}{4\gamma_{\mathrm{I}}}.$ (44)
Two independent solutions of the confluent-hypergeometric equation are $M(\mu;1+\nu;z):={}_{1}F_{1}(\mu;1+\nu;z)$ and $U(\mu;1+\nu;z)$ defined in Eq. (79) (see Appendix A for the definitions of them and their properties). Hence, the general solution of Eq. (40) on the region can be expressed as
@@ III.2 Solution of Schrödinger equation
First, we consider this situation. Because a solution should not diverge at $r\to 0$ and $r\to\infty$, one gets
@@
$\Phi_{\mathrm{I},\mathrm{M}}^{(1)}(r):=r^{\nu_{\mathrm{I}}}e^{-\frac{\gamma_{\mathrm{I}}r^{2}}{2}}M(\mu_{\mathrm{I}};1+\nu_{\mathrm{I}};\gamma_{\mathrm{I}}r^{2}),\quad\Phi_{\mathrm{I},\mathrm{U}}^{(1)}(r):=r^{\nu_{\mathrm{I}}}e^{-\frac{\gamma_{\mathrm{I}}r^{2}}{2}}U(\mu_{\mathrm{I}};1+\nu_{\mathrm{I}};\gamma_{\mathrm{I}}r^{2})$ (47)
@@
Here, we can rewrite Eq. (50) explicitly as
Interestingly, the solutions of Eq. (50) are classified into two types: trivial solutions with integer eigenvalues and nontrivial solutions. As will be seen, only the nontrivial solutions are relevant.
@@ IV.1.1 Eigenvalues and eigenfunctions
Figure 4 represents the values of $\det\mathcal{M}(\lambda)$ against $\lambda$ using Eqs. (50) and (53) for $T=1$, $k_{\mathrm{in}}=1.0$, $k_{\mathrm{mid}}=-0.5$, $k_{\mathrm{out}}=0.8$, $\xi=1.0$, and $\alpha=3.0$. It exhibits various solutions of $\det(\mathcal{M}(\lambda))=0$. Interestingly, odd crossing points with $\partial_{\lambda}\det(\mathcal{M}(\lambda))>0$ are expressed as integers, i.e., $\lambda=0,1,2,3,4,\cdots$. However, these zeros cannot be the eigenvalues of Eq. (40) because they satisfy another condition

### 01 Apr 2026 v3

@@ Abstract
We present an exactly solvable model of the Mpemba effect in an overdamped Langevin system confined to a two-dimensional, radially symmetric bistable potential. The potential is constructed as a piecewise quadratic-logarithmic function that is continuous and differentiable at the matching radii, enabling an exact mapping of the corresponding Fokker-Planck operator to a Schrödinger-type eigenvalue problem. The relaxation spectrum and eigenmodes are obtained analytically in each region in terms of confluent hypergeometric functions, with eigenvalues determined from matching conditions. Focusing on isotropic equilibrium initial states at inverse temperature $\beta_{\rm ini}$ quenched to a bath at inverse temperature $\beta$, we derive explicit expressions for the mode amplitudes governing long-time relaxation. We demonstrate that the coefficient of the slowest mode exhibits a non-monotonic dependence on $\beta_{\rm ini}$ and identify a sufficient crossing condition for the Kullback-Leibler divergence in terms of the two slowest modes, if the global minimum of the potential is located far away from the origin and the second minimum exists near the origin. For corresponding parameters, we demonstrate that the Mpemba effect can be realized. Our results provide a rare example of an analytically tractable two-dimensional model exhibiting anomalous relaxation without any confining walls, extending previous one-dimensional constructions with a hard wall and clarifying the role of radial geometry in nonequilibrium relaxation phenomena.
@@ I Introduction
The Mpemba effect has been reported in a broad spectrum of classical systems, including colloidal particles in optical traps [6, 7], thermostated granular fluids [5, 8, 9, 10, 11, 12, 13], optical resonators [14, 15, 16], inertial suspensions [17, 18], spin glasses [19], molecular binary mixtures [20], phase transitions [21, 22], and a wide variety of Markovian model [4, 23, 24, 25, 26, 27, 28], non-Markovian models [29, 30] and active matter [31]. In addition, quantum analogues of the effect have been demonstrated in open quantum systems [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], establishing the Mpemba effect as a general nonequilibrium relaxation phenomenon.
@@
So far, we have only results for the motion of a particle in a discrete-level model, piece-wise linear potentials, and square box potentials [4, 5, 6, 57, 27]. To the best of our knowledge, there are no analytical studies of the motion of a particle confined to a smooth bistable potential, even in one-dimensional (1D) settings. Experimentally and theoretically, one-dimensional systems differ fundamentally from higher-dimensional systems, such as two-dimensional (2D) systems. Quite recently, Liu et al. [53, 54] have analyzed the motion in 1D asymmetric bistable potentials and found that the metastable picture is not correct for the realization of the Mpemba effect, but a confining wall is necessary to observe the Mpemba effect. To understand the robustness of the Mpemba effect, it is natural to analyze a 2D system to clarify the differences and common features in 1D and 2D systems, including whether confining walls are necessary to observe the Mpemba effect even in 2D.
@@
The contents of this paper are as follows. In the next section, we explain the model and setup, including the introduction of a solvable potential, and the mapping onto the Schrödinger equation. In Sec. III shows the mode-analysis of the mapped Schrödinger equation to obtain the solution of Schrödinger equation. In Sec. IV, we demonstrate the occurrence of the Mpemba effect by using the two-mode approximation. In Sec. V, we discuss our results. In Sec. VI, we present the summary of our results. In Appendix A, we briefly summarize the properties of hypergeometric functions we use. In Appendix B, we describe how the eigenmodes are determined. In Appendix C, we present the condition of the crossing of the KL divergence. In Appendix D, we discuss the contribution of higher-order terms of mode analysis. In Appendix E, we present the detailed conditions to have a peak of the slowest eigenmode against the initial temperature, mathematically.
@@ II.1 Formulation
Figure 1: A schematic of a bistable radially symmetric potential.
Let us consider the motion of a particle at position $\bm{r}$ in a two-dimensional system. Here, the particle moves under the influence of a bistable potential $V(r)$ with $r:=|\bm{r}|$ (see Fig. 1). We assume that the motion of the particle obeys the dimensionless Langevin equation as
@@
with $T$ being the bath temperature.
@@
Note that $P(\bm{r},t)$ should be isotropic if the initial condition is isotropic, i.e. $P(\bm{r},t)=P(r,t)$ because the dynamics is governed by a radially symmetric potential $V(r)$.
@@ II.2 A choice of solvable potential
In general, it is difficult to obtain the exact solution of the Fokker-Planck equation with a double-well potential. However, for a suitably chosen form of the potential $V(r)$, we can solve the Fokker-Planck equation. We therefore construct an exactly solvable radially symmetric double-well potential $V(r)$ in two dimensions. The requirements are: 1) $V(r)$ is continuous and differentiable across its domains, 2) $V(r)$ has a minimum at $r=0$, a maximum at $r=\xi$, and a second minimum at $r=\alpha>\xi$.
@@
where $r_{\pm}$ denote matching points, and $C_{\rm in},C_{\rm mid}$, and $C_{\rm out}$ are, respectively, constants to ensure continuity. As will be shown, the conditions $b_{\mathrm{mid}}\propto\xi^{2}$ and $b_{\mathrm{out}}\propto\alpha^{2}$ should be satisfied. Notably, this specific form leads to an exactly solvable Fokker-Planck equation.
The derivatives in each region are given by
@@
The second derivative of $V(r)$ are given by
@@
The potential $V(r)$ in Eq. (10) has two minima at $r=0$ and $r=\alpha$. Then, the values of $V(r)$ at these minima are, respectively, given by $V(0)=0$, and
@@
where we have used Eqs. (9), (13), (14), and (15). We distinguish three cases: (i) $V(\alpha)<0$, (ii) $V(\alpha)=0$, and (iii) $V(\alpha)>0$. For the representative choice $k_{\mathrm{in}}=k_{\mathrm{out}}=\xi=1$, the marginal line for $V(\alpha)=0$ is shown in Fig. 2.
@@ II.3.1 Framework
Requiring that $V_{S}(r)$ has extrema at $r=0$, $\xi$, and $\alpha$ yields
@@
We note the following: i) Each segment of $V_{S}(r)$ is solvable quadratic plus inverse-square potential. ii) The solutions of the corresponding radial Schrödinger equation are expressible in terms of confluent hypergeometric functions. iii) The matching conditions at $r=r_{-}$ and $r=r_{+}$ enforce continuity of both $V(r)$ and $V^{\prime}(r)$, which automatically ensure the continuity of $V_{S}(r)$ and the physical wavefunctions. This construction provides a fully solvable model for a two-dimensional bistable potential with spherical symmetry.
@@
where $\varphi_{m,n}(r)$ is related to the left and right eigenfunctions as
@@
When we begin with an isotropic (equilibrium) condition Eq. (6), the solution of Schrödinger equation is independent of the second index $n$. Thus, the orthonormal relation
@@
$\langle\hat{A}\rangle(t)\approx 2\pi\int_{0}^{\infty}drrP_{\mathrm{eq}}(r,\beta)\hat{A}+2\pi\int_{0}^{\infty}drr\hat{A}\left[e^{-\beta V(r)/2}\left\{a_{2}\phi_{2}(r)e^{-\tilde{\lambda}_{2}t}+a_{3}\phi_{3}(r)e^{-\tilde{\lambda}_{3}t}\right\}\right],$ (34)
@@
Note that one-dimensional model can be solved using a set of harmonic and anti-harmonic potentials [54]. The idea of using harmonic and anti-harmonic potentials to describe tunneling in an asymmetric double-well potential in quantum mechanics was introduced by Dekker [60] and later developed by Song [61, 62]. Our analysis shares similarities with previous papers [60, 61, 62], the target is completely different from them.
@@ II.3.2 Condition to observe the Mpemba effect
Let us consider Eqs. (II.3.1) and (34) in the large $t$ regime, where the relaxation term proportional to $a_{2}$ is only relevant. Then, the previous papers [4, 51, 53, 54] have shown that the condition to observe the Mpemba effect is that $a_{2}$ has a maximum or a minimum against $\beta_{\mathrm{ini}}$. Indeed, this condition is necessary because the peak of $a_{2}$ at $\beta_{\mathrm{ini}}^{*}$ corresponds to the slowest initial temperature to approach the final equilibrium state after the quench. This suggests that two relaxation processes, one starting from $\beta_{\mathrm{ini}}^{*}$ and the other from $\beta_{\mathrm{ini}}<\beta_{\mathrm{ini}}^{*}$, may cause a crossing of a thermodynamic observable. However, this condition may be insufficient, as it does not guarantee that the slowest eigenmode crosses another relaxation mode proportional to $a_{3}$. To improve the defect of the previous condition, only $a_{2}$, we will propose a more direct crossing condition for a monotonic measure, such as the KL divergence, to observe the Mpemba effect later.
@@ II.3.3 Matching conditions
Since we adopt the connected potential Eq. (9), which can be converted into the effective potential Eq. (21), we should impose the connection conditions at $r=r_{-}$ and $r=r_{+}$. The connection conditions are given by
@@
where $P_{\mathrm{in}}(r;t)$, $P_{\mathrm{mid}}(r;t)$, and $P_{\mathrm{out}}(r;t)$ are $P(r,t)$ for $0\leq r<r_{-}$, $r_{-}\leq r<r_{+}$, and $r\geq r_{+}$, respectively. Similarly, $J_{r}(r,\theta;t):=\bm{e}_{r}\cdot\bm{J}(r,\theta;t)$ is the radial component of the current $\bm{J}$ introduced in Eq. (5), and $J_{r,\mathrm{in}}$, $J_{r,\mathrm{mid}}$, and $J_{r,\mathrm{out}}$ are $J_{r}$ for $0\leq r<r_{-}$, $r_{-}\leq r<r_{+}$, and $r\geq r_{+}$, respectively. Since $V^{\prime}(r)$ is continuous at the connection points, the continuity condition of the current reduces to the continuity condition of $\partial P(r,t)/\partial r$ if the radial current $J_{r}$ is regular at the connection points. Using Eqs. (5) and (II.3.1), the continuity conditions can read
@@ III.1 Eigenvalue equation
Let us rewrite $\mathbb{H}$ introduced in Eq. (19) in polar coordinates, $r:=\sqrt{x^{2}+y^{2}}$, under the assumption of circular symmetry:
@@
$\mu_{\mathrm{I},n}:=\frac{1+\nu_{\mathrm{I}}}{2}-\frac{\beta\left(\lambda_{n}-k_{\mathrm{I}}-\tfrac{k_{\mathrm{I}}|b_{\mathrm{I}}|}{2T}\right)}{4\gamma_{\mathrm{I}}}.$ (44)
Two independent solutions of the confluent-hypergeometric equation are $M(\mu;1+\nu;z):={}_{1}F_{1}(\mu;1+\nu;z)$ and $U(\mu;1+\nu;z)$ defined in Eq. (78) (see Appendix A for the definitions of them and their properties). Hence, the general solution of Eq. (40) in the region can be expressed as
@@ III.2 Solution of Schrödinger equation
First, we consider this situation. Because a solution should not diverge at $r\to 0$ and $r\to\infty$, it follows that
@@
$\Phi_{\mathrm{I},\mathrm{M}}^{(1)}(r):=r^{\nu_{\mathrm{I}}}e^{-\tfrac{\gamma_{\mathrm{I}}r^{2}}{2}}M(\mu_{\mathrm{I}};1+\nu_{\mathrm{I}};\gamma_{\mathrm{I}}r^{2}),\quad\Phi_{\mathrm{I},\mathrm{U}}^{(1)}(r):=r^{\nu_{\mathrm{I}}}e^{-\tfrac{\gamma_{\mathrm{I}}r^{2}}{2}}U(\mu_{\mathrm{I}};1+\nu_{\mathrm{I}};\gamma_{\mathrm{I}}r^{2})$ (47)
@@
Interestingly, the solutions of Eq. (50) are classified into two types: trivial solutions with integer eigenvalues and nontrivial solutions. As will be seen, only the nontrivial solutions are relevant. For the explicit procedure to determine the eigenmodes, see Appendix B.
@@ IV.1.1 Eigenvalues and eigenfunctions
Figure 4 represents the values of $\det\mathcal{M}(\lambda)$ against $\lambda$ using Eqs. (50) and (52) for $T=1$, $k_{\mathrm{in}}=1.0$, $k_{\mathrm{mid}}=-0.5$, $k_{\mathrm{out}}=0.8$, $\xi=1.0$, and $\alpha=3.0$. It exhibits multiple solutions of $\det(\mathcal{M}(\lambda))=0$. Interestingly, odd crossing points with $\partial_{\lambda}\det(\mathcal{M}(\lambda))>0$ are expressed as integers, i.e., $\lambda=0,1,2,3,4,\cdots$. However, these zeros cannot be the eigenvalues of Eq. (40) because they satisfy another condition